OpenStudy (anonymous):

Movie Preference Action Drama Comedy Row totals High School 400 230 120 750 College 200 510 280 990 Column totals 600 740 400 1740 Julie mistakenly calculated the conditional relative frequency for college students who like comedy as being 70%. What did Julie actually calculate, and what is the correct answer?

OpenStudy (anonymous):

She calculated the joint relative frequency for college/comedy. The correct value of conditional relative frequency for college students who like comedy is 28%. She calculated the conditional relative frequency for comedy films that are liked by college students. The correct value of conditional relative frequency for college students who like comedy is16%. She calculated the conditional relative frequency for comedy films that are liked by college students. The correct answer is 28%. She calculated the joint relative frequency for college/comedy. The correct answer is 16%.

OpenStudy (misssunshinexxoxo):

Which do you believe it is?

OpenStudy (nincompoop):

this can be solved without guessing

OpenStudy (nincompoop):

If I had 10 marbles, 70% of it would be 7. If I had 40 wives, 50% of it would be 25. Notice that in order to get the value of a certain percentage, the total must be included. So if you had a total sample size of 1000, then 70% of it is 700. But what if I am actually looking at the 300 and not the 7000, then you would know what percentage it is (since per cent pertains to per 100) so 300. This can get more complicated, but it would be easy if you know how to solve it. \(\huge \frac{700}{1000} \times 100 \rightarrow \frac{group ~ of~ people~to~count}{total }\times 100\)